# PhoenixOS: Concurrent OS-level GPU Checkpoint and Restore with Validated Speculation

### v2

@@ Introduction
Figure 11 (a) measures the checkpoint overhead, with the checkpoints done at the beginning of the iteration. The stall is calculated by first subtracting the total training time with and without the checkpoint and then normalized to a single iteration time. We can see that PhOS reduced the checkpoint overhead by 70–160% compared to Singularity thanks to the concurrent checkpoint. Notably, on Llama2-13B training, the overhead of PhOS is only 185 ms, while Singularity is 3.2 s, bottlenecked by transferring the checkpoint data (72 GB) to the host memory (32 GBps PCIe). For reference, the iteration time is 6.9 s.
Figure 12 further compares the average wasted GPU hours with the optimal checkpoint frequency, whose formula is listed in §A.1, the same as a priori work [28]. Since the wasted time is related to the failure frequency, we set a failure ratio of one GPU fails per hour reported by various industry reports [78, 46]. The results show that PhOS saves up to 22–86% GPU hours compared with other approaches. This is because PhOS enables a higher checkpoint frequency to save wasted GPU hours due to recomputation. For example, in Llama2-13B training, PhOS enables a checkpoint frequency of 279 per hour, while the optimal setup for Singularity is 67.
@@
Live migration: evaluated metric and results.   To evaluate the effectiveness of PhOS in live migration scenarios, we measure the downtime when migrating applications between nodes. Figure 13 shows the timeline of application performance during migration. The results show that PhOS incurs a minimal downtime thanks to the concurrent execution capabilities. Notably, PhOS only incurs 3.3 and 3.7 seconds of downtime for migrating Llama2-13B training and LLama2-70B inference, respectively. In comparison, Singularity incurs 10.2 and 12.35 seconds of downtime, which is dominated by copying data from the source GPU to the target. In Llama2-13B training, copying data through 100 Gbps RDMA takes at least 9.8 seconds.
@@
Figure 14: The serverless function execution time under coldstart. The Exe. states for the function execution time. All the workloads are inference.
@@
Figure 15: An analysis of the runtime validator overhead on (a) training and (b) inference workloads. (c) Analysis of the ratios of the kernels instrumented with the validator. -T is abbreviated for training while -I is for inference.
Serverless function startup: evaluated metric and results.   We evaluate the effectiveness of PhOS in GPU serverless workloads by measuring the end-to-end execution time, i.e., considering both startup and application function execution time [5, 20]. We stored the function checkpoint in host DRAM and measured the execution time under cold start with OS-level restore. We choose inference workloads because training workloads are stable and do not need serverless-style startup. Figure 14 shows the results: PhOS has the fastest execution time: on average, it improves upon cuda-checkpoint and Singularity by 24$\times$ and 16$\times$ respectively, thanks to the eliminated GPU context creation time as well as its ability to overlap copying data stalls with concurrent restore.
@@
Figure 16: A breakdown of the checkpoint overhead on (a) PhOS CoW, (b) without our prioritized PCIe transfer optimization and (c) Singularity. Workload: Llama2-13B Training. Note that we omitted the concurrent data copy of PhOS.
@@
Breakdown of the copy-on-write protocol.   Figure 16 shows the breakdown of the checkpoint stall caused by different checkpoint methods on a Llama2-13B training workload. Other workloads share a similar trend. All the system checkpoints at the beginning of the iteration: the optimal checkpoint timing, as analyzed in §8.3. We can see that first, though the quiescing phase will stop the application execution, its absolute time (10 ms) is extremely small so this phase has a negligible overhead. The absolute time is small because (1) ongoing kernels are short-lived and (2) coordinating between threads with RDMA to reach a global quiesce is extremely efficient. Second, though PhOS may stall application due to copy-on-write, the aggregated stalls are small as long as PhOS can timely finish the data copy.
@@
Effectiveness of the prioritized application PCIe transfer. We further evaluate the effectiveness of our prioritized application PCIe transfer optimization described in §5. As shown in Figure 16, without this optimization (b), applications still suffers stalls due to PCIe bandwidth competition, even when they have not written data that may cause an incorrect checkpoint (a). This is because the GPU kernels are waiting to load the training set from the CPU, which is blocked by the checkpoint data copy due to the limited number of GPU DMA engines on the GPUs.
@@
Figure 17: A breakdown of the checkpoint overhead on (a) PhOS recopy, (b) without our coordinated CPU GPU checkpoint optimization and (c) Singularity. Workload: Llama2-70B Inference. Note that we omitted the concurrent data copy of PhOS.
@@
Breakdown of the recopy protocol. Figure 17 shows the breakdown of executing our recopy protocol on a Llama2-70B inference workload. We can see that the downtime of PhOS is dominated by the recopy time. The downtime is longer for recopy when compared with CoW because it not only ensures that the checkpoint is correct, but it is also fresh. Nevertheless, the downtime is still orders of magnitude smaller than a stop-the-world approach (Singularity) because the recopied data is orders of magnitude smaller (2.1 vs. 9.7 seconds).
@@
Effectiveness of the coordinated CPU and GPU checkpoint. Figure 17 further evaluates the effectiveness of our coordinated CPU and GPU checkpoint optimization in §5, aiming to reduce the amount of data for the recopy phase. Thanks to the optimized GPU concurrent copy timing, PhOS enjoys 47% smaller recopy time than without optimization (b). The recopied data reduces from 50 to 27 GB per GPU. Since the copy is a bulk load procedure, the reduced transferred size directly translates to a smaller recopy time.
@@
Figure 18: A breakdown of the restore on (a) PhOS concurrent restore, and (b) Singularity. Workload: Llama2-13B inference. Note that we omitted the concurrent data copy of PhOS. R stall states for restore stall time.
@@
Breakdown of the concurrent restore. Figure 18 shows the breakdown of concurrent restoring a Llama2-13B inference process using PhOS. We can see that compared to the stop-the-world restore (b), the improvements of PhOS come from two factors: (1) the eliminated context creation and (2) overlapping the data copy with the kernel execution. Note that during application execution, PhOS still copies the data in the background: this perfectly aligns with how inference processes access data nowadays: before executing the first layer of inference, we can concurrently restore the second layer’s data. Unlike previous work [75, 6], PhOS achieves so transparently.
@@
The runtime overhead of tracing GPU read and write set. Figure 15 (a) and (b) report the runtime overhead of runtime validator on different applications. We observe a relatively small slowdown of 1–12% for various workloads. The observable overhead is small because: (1) the additional check only happens when the kernel accesses the global memory, which is less frequent for high computational efficiency; and (2) the instrumented kernels are only a small portion of the overall kernels in the workloads, as shown in Figure 15 (c).
@@
Figure 19: Impact of checkpoint timing on the performance of Llama2-13B training. The blank in App is the stalled caused by the CoW. The red indicates the data copy time. Note that we omitted the concurrent data copy of PhOS.
@@
Figure 20: A heatmap showing the read and write sets traced by our speculated validation method on a Llama2-13B training workload.
@@
Table 3: A study of the success rate of our speculation-based GPU read and write sets tracing beyond our evaluated applications.
@@
Rodinia [10] 44 / 1 48,610 / 20
Parboil [65] 18 / 0 43,473 / 0
vLLM [35] 66 / 0 13,625 / 0
TVM [11] 607 / 0 186,244 / 0
FlashInfer [3] 69 / 0 15,265/ 0
For all the common GPU applications evaluated by PhOS, we don’t meet any speculation failure. To evaluate the feasibility of our speculation, we conduct a comprehensive study across diverse GPU applications including supercomputing workloads (Rodinia [10] and Parboil [65]), dynamically generated kernels from AI compilers (TVM [11]), and highly optimized handwritten kernels (vLLM [35] and FlashInfer [3]) in Table 3. To evaluate the success rate of our speculation, we run typical tasks provided by them (we term each task instance). For Rodinia and Parboil, we ran all benchmarks except for ones that had bugs or used outdated CUDA APIs (e.g., dwt2d). For TVM, we ran inference on their available models other than our previous evaluated (e.g., DenseNet, YOLOv4). Out of all instances, only one kernel (from Rodinia) failed our speculation—a fairly dated supercomputing application. The failure occurred because it reads a buffer pointed to by a global variable not listed in the arguments.
@@ Discussion
Supporting advanced GPU features like CUDA graph.   CUDA graph [4] allows the CPU to submit a batch of kernels. PhOS supports tracing reads and writes of kernels with CUDA graph, inspired by Medusa [77]. Specifically, CUDA graphs can be utilized as follows: (1) explicitly invoking APIs such as cudaGraphAddKernelNode, or (2) implicitly constructing them through the driver using cudaStreamBeginCapture. Both methods require explicit driver API calls, so they are compatible with our speculative approach.
GPU hardware extensions and OS-level ABI for concurrent C/R.   PhOS is the first system to show the effectiveness of concurrent GPU checkpoint and restore as well as how to realize it in major GPUs. While hardware extensions like GPU dirty bit could simplify and accelerate PhOS’s implementation, they are less flexible, e.g., a hardware dirty bit alone cannot support our other protocols like soft copy-on-write. On the other hand, an OS-level ABI enabling developers to specify kernel buffer access patterns would fully utilize PhOS’s capabilities. Yet modifying numerous existing custom kernels to adopt this ABI is non-trivial. How to develop systematic fine-grained kernel memory access tracing methods beyond our speculative approach is left as our future work.
@@
Checkpoint and restore.  PhOS continues the line of research on C/R, with a focus on efficient OS-level GPU C/R [21, 41, 29, 62, 66, 18, 36, 30, 73, 64], a feature currently only available on CPU. A recent work (whose submission time is 2 months after our first released preprint [^2] https://arxiv.org/abs/2405.12079)—gCROP [75], uses concurrent restore to accelerate process startup (only) on AMD GPUs. They ensure concurrent restore correctness with significant application modifications, which PhOS achieves the same effect without application modifications using our novel validated speculation. They cannot support multiple GPUs, their approach is infeasible on NVIDIA GPUs, and most importantly, they don’t support concurrent checkpoint—though their paper title misleadingly claims so. GPU snapshot [37] realizes dirty bit with simulated GPU hardware extension, which to the best of our knowledge, no real hardware implementation exists. PhOS is the only system that supports both concurrent OS-level checkpoint and restore for GPU applications on multi-GPU and major GPU vendors (i.e., NVIDIA GPUs).
Analyzing GPU kernels.   Many existing works analyze GPU programs (kernels), either statically or during runtime [12, 39, 8, 9, 34, 38, 40, 42]. PhOS takes a different approach: instead of analyzing the program, we analyze the kernel arguments for speculation. This avoids the incompleteness inherent in GPU kernel analysis (as GPU kernels are Turing complete) and works well for concurrent GPU checkpoint and restore.
@@ Acknowledgment
We would like to thank the reviewers from SOSP’24, OSDI’24 and SOSP’25 for their insightful feedback. We sincerely thank Zhiyuan Dong, Dingyan Zhang and Xiating Xie for giving us feedback on the early draft of this paper. We would also like to thank Shaoxun Zeng for sharing details on how PhOS can be adapted to support CUDA graph. This work was supported in part by the National Key Research & Development Program of China (No. 2022YFB4500700), the Fundamental Research Funds for the Central Universities, the National Natural Science Foundation of China (No. 62202291, 62272291), as well as a research grant from Huawei.
@@ Appendix A Appendix
Table 4: Detailed setups of applications evaluated in §8. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama-3-70B training.
ResNET-${}^{\text{152M}}$) PPO-${}^{\text{336M}}$) Stable Diffusion-${}^{\text{1B}}$) Llama2-${}^{\text{13B}}$) Llama3.3-${}^{\text{70B}}$)
@@
Library Torchvision [60] OpenAI Gym [57] HuggingFace [24] Meta Llama 2 [43] Meta Llama 3.3 [44]
@@
[…]
Solving the optimal $f$. For a given job, the distribution of failures ($F$), the checkpoint overhead ($O$), and the restore time ($R$) are static and can be profiled online [46], so we only need to solve the checkpoint frequency $f$ to minimize the total GPU hours wasted. By differentiating the total GPU hours wasted with respect to $f$, we got:
@@
Figure 21: An illustration of how applications can use PhOS SDK to control the checkpoint timing. Lines with + are code interacting with PhOS.
@@
Table 4 shows the detailed buffer allocation information as well as the number of GPU kernels information of our evaluated applications. The detailed setups of training workloads are as follows: For ResNet50 training, we use the CIFAR-10 dataset and configure a batch size of 32 for training. For PPO training, we use the gym code [57]. For Stable Diffusion training, we use the SD v1-4 model and a batch size of 1,536 for each GPU. Finally, for Llama training, we use a distributed configuration of 8TP, the batch size is 4, restricted by the available GPU memory. All the training workloads use the AdamW optimizer.

### v3

@@ Introduction
Figure 11 (a) measures the checkpoint overhead, with the checkpoints done at the beginning of each training iteration. The stall is calculated by first subtracting the total training time with and without the checkpoint and then normalized to a single iteration time. We can see that PhOS reduced the checkpoint overhead by 70–160% compared to Singularity thanks to the concurrent checkpoint. Notably, on Llama2-13B training, the overhead of PhOS is only 185 ms, while Singularity is 3.2 s, bottlenecked by transferring the GPU checkpoint data (72 GB) to the host memory (via 32 GBps PCIe). For reference, the iteration time is 6.9 s.
Figure 12 further compares the average wasted GPU hours caused by different checkpoint methods. Since the wasted time is related to the failure frequency and the checkpoint frequency, we set a failure ratio of one GPU failure per hour reported by various industry reports (DBLP:journals/corr/abs-2205-01068, ; DBLP:conf/fast/MohanPC21, ), and then calculate the optimal checkpoint frequency for each system according to the formula in §A.1. Note that different systems require distinct checkpoint frequencies: PhOS achieves optimal performance with 279 checkpoints per hour, whereas Singularity performs best at 67. PhOS can set a higher frequency thanks to the reduced checkpoint overhead. The results show that PhOS saves up to 22–86% GPU hours compared with other approaches: the higher checkpoint frequency of PhOS saves wasted GPU hours due to recomputation, while our concurrent checkpoint minimizes wasted GPU time caused by high checkpoint overhead with high checkpoint frequency. Interestingly, a well-implemented OS-level GPU C/R even outperforms a simple user-level checkpointing (torch.checkpoint) that uses PyTorch’s save.
@@
Live migration: evaluated metric and results.    To evaluate the effectiveness of PhOS in live migration scenarios, we measure the downtime when migrating applications between nodes with different systems. Figure 13 shows the timeline of application performance during migration. We can see that PhOS incurs a minimal downtime thanks to the concurrent execution capabilities. Notably, PhOS only incurs 3.3 and 3.7 seconds of downtime for migrating Llama2-13B training and LLama2-70B inference, respectively. In comparison, Singularity incurs 10.2 and 12.35 seconds of downtime, which is dominated by copying data from the source GPU to the target. In Llama2-13B training, copying data through 100 Gbps RDMA takes at least 9.8 seconds.
@@
Figure 14. The serverless function execution time under coldstart. The Exe. states for the function execution time. All the workloads are inference.
@@
Figure 15. An analysis of the runtime validator overhead on (a) training and (b) inference workloads. (c) Analysis of the ratios of the kernels instrumented with the validator. -T is abbreviated for training while -I is for inference.
[…]
@@
Figure 16. A breakdown of the checkpoint overhead on (a) PhOS CoW, (b) without prioritized PCIe transfer optimization and (c) Singularity. Workload: Llama2-13B Training. Note that we omitted the concurrent data copy of PhOS.
@@
Breakdown of the copy-on-write protocol.    Figure 16 shows the breakdown of the checkpoint stall caused by different checkpoint methods on a Llama2-13B training workload. Other workloads share a similar trend. All the system checkpoints at the beginning of the iteration: the optimal checkpoint timing, as analyzed later in §8.4. We can see that first, though the quiescing phase will stop the application execution, its absolute time (100 ms) is much smaller so this phase has a negligible overhead. The absolute time is small because (1) ongoing kernels are short-lived and (2) coordinating between threads with RDMA to reach a global quiesce is extremely efficient. Second, though PhOS may stall application due to copy-on-write, the aggregated stalls are small.
@@
Effectiveness of the prioritized application PCIe transfer.    As shown in Figure 16, without our prioritized application PCIe transfer optimization described in §5 (b), applications still suffers stalls due to PCIe bandwidth competition. This is because the GPU kernels are waiting to load the training set from the CPU, which is blocked by the checkpoint data copy due to the limited number of GPU DMA engines on the GPUs.
@@
Figure 17. A breakdown of the checkpoint overhead on (a) PhOS recopy, (b) without coordinated CPU GPU checkpoint and (c) Singularity. Workload: Llama2-70B Inference. Note that we omitted the concurrent data copy of PhOS.
@@
Breakdown of the recopy protocol.    Figure 17 shows the breakdown of executing our recopy protocol on a Llama2-70B inference workload. We can see that the downtime of PhOS is dominated by the recopy time. Compared to CoW, the recopy downtime is longer because it not only ensures that the checkpoint is correct, but it is also fresh. Nevertheless, the downtime is still orders of magnitude smaller than a stop-the-world approach (Singularity) because the recopied data is much smaller (2.1 vs. 9.7 seconds).
@@
Effectiveness of the coordinated CPU and GPU checkpoint.    Figure 17 further evaluates the effectiveness of our coordinated CPU and GPU checkpoint optimization in §5, aiming to reduce the amount of dirty data for the recopy phase. Thanks to the optimized GPU concurrent copy timing, PhOS enjoys 47% smaller recopy time than without optimization (b). The recopied data reduces from 50 to 27 GB per GPU. Since the copy is a bulk load procedure, the reduced transferred size directly translates to a reduced recopy time.
@@
Figure 18. A breakdown of the restore on (a) PhOS concurrent restore, and (b) Singularity. Workload: Llama2-13B inference. Note that we omitted the concurrent data copy of PhOS. R stall states for restore stall time.
@@
Breakdown of the concurrent restore.    Figure 18 shows the breakdown of concurrent restoring a Llama2-13B inference process using PhOS. We can see that compared to the stop-the-world restore (b), the improvements of PhOS come from two factors: (1) the eliminated context creation and (2) overlapping the data copy with the kernel execution. Note that during application execution, PhOS still copies the data in the background: this perfectly aligns with how inference processes access data nowadays: before executing the first layer of inference, we can concurrently restore the second layer’s data. Unlike previous work (10.1145/3698038.3698510, ; pipeswitch, ), PhOS achieves so transparently.
@@
The runtime overhead of tracing GPU read and write set.    Figure 15 (a) and (b) report the runtime overhead of the runtime validator on different applications. We observe a relatively small slowdown of 1–12% for various workloads. The observable overhead is small because: (1) the additional check only happens when the kernel accesses the global memory, which is less frequent for high computational efficiency; and (2) the instrumented kernels are only a small portion of the overall kernels in the workloads. Figure 15 (c) reports the number of instrumented kernels invoked during the concurrent checkpoint and restore: We can see that most of the kernels are not instrumented, e.g., in Llama2-13B inference, only 12% of the invoked kernels are instrumented.
Impact of on-device memory used on performance
Figure 19. An analysis of the copy-on-write overhead with different on-device buffer sizes.
Figure 19 shows the impact of the on-device buffer size on the checkpoint overhead when checkpointing a Llama2-13B training workload. We choose this workload because it has the largest memory footprint such that little room is left for the on-device buffer (see Table 4). We report the increased stall time caused by the checkpoint, where the time is normalized to the stall time with our maximum configurable on-device buffer size (2 GB). We can see that checkpoint overhead generally decreases with the on-device buffer size, but the difference is small, e.g., even without on-device buffer, PhOS incurs up to 9.1 % more stalls. This is because not all copy-on-write operations require stalling the kernel, so the slow copy through PCIe with insufficient on-device buffer can be hidden.
@@
Figure 20. Impact of checkpoint timing on the performance of Llama2-13B training. The blank in App is the stalled caused by the CoW. The red indicates the data copy time. Note that we omitted the concurrent data copy of PhOS.
@@
Figure 21. A heatmap showing the read and write sets traced by our speculated validation method on a Llama2-13B training workload.
@@
Table 3. A study of the success rate of our speculation-based GPU read and write sets tracing beyond our evaluated applications.
@@
Rodinia (DBLP:conf/iiswc/CheBMTSLS09, ) 44 / 1 48,610 / 20
Parboil (stratton2012parboil, ) 18 / 0 43,473 / 0
vLLM (vllm, ) 66 / 0 13,625 / 0
TVM (DBLP:conf/osdi/ChenMJZYSCWHCGK18, ) 607 / 0 186,244 / 0
FlashInfer (flashinfer, ) 69 / 0 15,265/ 0
For all the common GPU applications evaluated by PhOS, we don’t meet any speculation failure. To evaluate the feasibility of our approach, we conduct a comprehensive study across diverse GPU applications including supercomputing workloads (Rodinia (DBLP:conf/iiswc/CheBMTSLS09, ) and Parboil (stratton2012parboil, )), dynamically generated kernels from AI compilers (TVM (DBLP:conf/osdi/ChenMJZYSCWHCGK18, )), and highly optimized handwritten kernels (vLLM (vllm, ) and FlashInfer (flashinfer, )) in Table 3. To evaluate the success rate of our speculation, we run typical tasks provided by them (we term each run of a task instance). For Rodinia and Parboil, we ran all benchmarks except for ones that had bugs or used outdated CUDA APIs (e.g., dwt2d). For TVM, we ran inference on their available models other than our previous evaluated ones (e.g., DenseNet, YOLOv4). Out of all instances, only one kernel (from Rodinia) failed our speculation—a fairly dated supercomputing application. The failure occurred because a kernel reads a buffer referenced by a global variable not listed in the launch arguments.
@@ Discussion
Supporting advanced GPU features.    CUDA graph (cuda-graph, ) is an advanced GPU feature that allows the CPU to submit a batch of kernels. PhOS supports tracing reads and writes of kernels within CUDA graph, inspired by Medusa (DBLP:conf/asplos/ZengXGCL25, ). Specifically, CUDA graphs can be utilized as follows: (1) explicitly invoking APIs such as cudaGraphAddKernelNode, or (2) implicitly constructing them through the driver using cudaStreamBeginCapture. Both methods require explicit driver API calls, so they are compatible with our API-hooking-based tracing.
GPU hardware extensions.  PhOS is the first to show the effectiveness of concurrent GPU checkpoint and restore as well as how to realize it in major GPUs. While hardware extensions like GPU dirty bit could simplify and accelerate PhOS’s implementation, relying on hardware extensions is less flexible, e.g., a hardware dirty bit alone cannot support our other protocols like soft copy-on-write.
Application-provided kernel access information.     Providing the kernel’s memory access to PhOS via developer annotation or just-in-time (JIT) compilation would fully utilize PhOS’s capabilities. Yet modifying numerous existing custom kernels for the annotation, as well as designing an effective JIT for checkpointing, is non-trivial. How to develop kernel memory access tracing methods beyond our speculative approach is left as our future work.
@@
Checkpoint and restore.  PhOS advances the research on checkpoint and restore (C/R), specifically focusing on efficient OS-level GPU C/R (egwutuoha2013survey, ; litzkow1997checkpoint, ; keykos, ; eros, ; Aurora, ; criu, ; laadan2010linux, ; hargrove2006berkeley, ; DBLP:conf/sosp/Wu0M023, ; DBLP:journals/corr/abs-2502-16631, ), a feature previously available only on CPUs. GPU snapshot (DBLP:conf/ics/LeeSHTKE19, ) realizes dirty bits through simulated GPU hardware extensions, which to the best of our knowledge, no actual hardware implementation currently exists. gCROP (10.1145/3698038.3698510, ) uses concurrent restore to accelerate process startup (only) on AMD GPUs, but requires significant application modifications to ensures correctness. Their approach is infeasible on closed-source GPUs (e.g., NVIDIA GPUs), does not support multi-GPU applications, and most importantly, lacks support for concurrent GPU checkpoint. In contrast, our work enables both concurrent GPU checkpoint and restore without application modifications through our novel validated speculation approach. PhOS is the first system to support both concurrent OS-level checkpoint and restore for GPU applications across multiple GPUs and major GPU vendors like NVIDIA GPUs.
Analyzing GPU kernels.    Many existing works analyze GPU programs (kernels), either statically or during runtime (DBLP:conf/nfm/ChiangGLR13, ; DBLP:conf/sigsoft/LiG10, ; DBLP:journals/toplas/BettsCDKQTW15, ; DBLP:conf/oopsla/BettsCDQT12, ; DBLP:conf/sosp/KamathB21, ; DBLP:conf/pldi/LeungGAGJL12, ; DBLP:conf/ppopp/LiLSGGR12, ; Mai2023HONEYCOMB, ). PhOS takes a different approach: instead of analyzing the program, we analyze the kernel arguments for speculation. This avoids the incompleteness inherent in GPU kernel analysis (as GPU kernels are Turing complete) and works well for concurrent GPU checkpoint and restore.
@@ Acknowledgment
We would like to thank the reviewers from SOSP’24, OSDI’24 and SOSP’25 for their insightful feedback. We sincerely thank Zhiyuan Dong, Dingyan Zhang and Xiating Xie for giving us feedback on the early draft of this paper. We would also like to thank Shaoxun Zeng for sharing details on how PhOS can be adapted to support CUDA graph. This work was supported in part by the National Key Research & Development Program of China (No. 2022YFB4500700), the Fundamental Research Funds for the Central Universities, the National Natural Science Foundation of China (No. 62202291, 62272291), as well as a research grant from Huawei. Corresponding author: Rong Chen (rongchen@sjtu.edu.cn).
@@ Appendix A Appendix
Table 4. Detailed setups of applications evaluated in §8. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama3.3-70B training.
[…]
@@
Library Torchvision (torchvision, ) OpenAI Gym (gym, ) HuggingFace (huggingface, ) Meta Llama 2 (llama, ) Meta Llama 3.3 (llama3, )
@@
$\text{Total GPU-hours wasted}\;=\;N\,F\,T\Bigl{(}R\;+\;\tfrac{N}{2\,f}\Bigr{)}\;+\;N\,O\,f\,T.$
Solving the optimal $f$.  For a given job, the distribution of failures ($F$), the checkpoint overhead ($O$), and the restore time ($R$) are static and can be profiled online (DBLP:conf/fast/MohanPC21, ), so we only need to solve the checkpoint frequency $f$ to minimize the total GPU hours wasted. By differentiating the total GPU hours wasted with respect to $f$, we got:
@@
Figure 22. An illustration of how applications can use PhOS SDK to control the checkpoint timing. Lines with + are code interacting with PhOS.
@@
Table 4 shows the detailed buffer allocation information as well as the number of GPU kernels information of our evaluated applications. The detailed setups of training workloads are as follows: For ResNet50 training, we use the CIFAR-10 dataset and configure a batch size of 32 for training. For PPO training, we use the gym code (gym, ). For Stable Diffusion training, we use the SD v1-4 model and a batch size of 1,536 for each GPU. Finally, for Llama training, we use a distributed configuration of 8TP (Tensor Parallelism) and 1DP (Data Parallelism), the batch size is 4, restricted by the available GPU resources. All the training workloads use the AdamW optimizer.